# GranQ: Granular Zero-Shot Quantization with Unified Layer-Channel Awareness

### v3

@@ Abstract
Zero-shot quantization (ZSQ) enables neural network compression without training data, which is crucial in restricted data access environments. However, existing ZSQ methods suffer from significant activation loss in low-bit environments owing to their coarse-grained scaling strategy. To address this issue, we propose GranQ, a novel ZSQ approach that leverages layer-channel awareness to minimize the quantization error. Unlike conventional layer- or channel-wise quantization, GranQ dynamically adjusts quantization granularity by considering both layer- and channel-level activation distributions. This enables fine-grained quantization while minimizing activation distortion. Additionally, we introduce vectorized activation quantization, which enables efficient parallel computation and reduces computational overhead while preserving accuracy. GranQ achieves superior performance compared with those of state-of-the-art ZSQ methods that employ quantization-aware training. With these findings, we anticipate that GranQ will inspire novel research directions beyond conventional ZSQ approaches focused on data generation and model training. The official code will be released upon completion of the review.
@@ 1 Introduction
Figure 1: Comparison of quantization strategies. Existing methods apply either per-tensor scaling with layer-wise quantization or per-channel scaling. GranQ combines per-channel scaling with per-layer iteration for more granular quantization.
@@
Neural network compression techniques have been extensively studied to make large-scale deep learning (DL) models practically deployable. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression [7, 12] include quantization [24, 30, 17], pruning [28, 16, 33, 6], knowledge distillation [21, 20, 40], and neural architecture search [45, 13]. Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model while maintaining its performance. However, a major limitation of quantization is the need for fine-tuning or calibration to maintain the performance of the full-precision (FP) model [17]. To address this limitation, zero-shot quantization (ZSQ) (i.e., data-free quantization) has been introduced [34, 41], which allows compressed models to maintain their performance without requiring access to the original training data. Since the introduction of ZeroQ [4], studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
@@
First, data generation studies focus on generating high-quality data to effectively train the Q model. A common approach uses batch normalization statistics from the FP model [41]. Recent studies focused on generating data adaptable to various bit-widths [36] and using diverse generative models such as stable diffusion guided by text prompts [32]. Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [18, 2]. Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [10, 14, 23, 26].
However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved. To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning. Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods. Notably, we discovered that layer-wise (per-tensor) quantization fails to accurately preserve activations in ZSQ, leading to coarse and inaccurate representations.
Based on this analysis, we introduce GranQ, a novel ZSQ approach. GranQ independently handles large activations in specific layers and channels by narrowing the range between the maximum and minimum values within each channel (Figure 1). This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 2. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, we implement vectorized operations for granular quantization with minimal computational overhead, enabling efficient quantization on GPUs. Our contributions can be summarized as follows:
  • We perform a comprehensive analysis of quantization errors in low-bit ZSQ settings. Our findings reveal that conventional activation quantization methods (e.g., layer- and channel-wise) lead to significant activation loss. This issue stems from coarse-grained activation quantization, which fails to preserve fine-grained details.
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. We introduce a fine-grained quantization mechanism, which considers every channel in each layer, to reduce activation loss. Additionally, we mitigate computational overhead using vectorized activation quantization. To the best of our knowledge, our approach is the first to address the ZSQ problem.
(a) Layer Quantization
(b) GranQ (Ours)
Figure 2: Comparison between (a) layer-wise quantization and (b) the proposed GranQ method on the CIFAR-10 dataset from AdaDFQ[36]. We visualize the activation of the first layer in ResNet-20 for the first image of the first batch. $w$ and $a$ denote weight and activation, respectively. Each sub-figure presents the activation graph for 32-bit FP (left) and 3-bit quantization (right). Although layer-wise quantization significantly distorts the original activation after quantization, GranQ effectively preserves it with minimal distortion.
2 Related Work
2.1 Quantization
Quantization can reduce the representational range of pre-trained deep neural networks (DNNs) and minimize memory usage. Quantization methods are categorized into PTQ and QAT based on the timing of the range reduction.
PTQ quantizes fully trained DNNs without any further training [17, 3], offering low computational cost and simple implementation. However, because PTQ directly applies quantization to the FP model, it struggles to properly adjust the range during quantization. To overcome this, calibration is used to select optimal quantization ranges by determining scaling factors. In contrast, QAT fine-tunes the Q model with quantization operations integrated into the training process, enabling direct optimization of activations [17]. In particular, it employs the straight-through estimator [42] to estimate gradients.
However, both PTQ and QAT have the limitation of requiring access to data. PTQ needs data for calibration, whereas QAT requires it for fine-tuning. In practice, obtaining training data is often challenging. To address this problem, ZSQ has been proposed, which performs quantization without using original data [41].
@@ 2.2 Zero-shot Quantization
As summarized in Table 1, ZSQ was initially introduced by ZeroQ [4], leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data [25, 32] and developing more effective methods to utilize them [10, 23, 26]. However, these approaches are still limited in preserving activations in low-bit settings. Therefore, a new quantization method is required to effectively minimize activation loss. In this study, we propose a novel approach to address this challenge.
@@ 3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the supplementary material.
@@ 3.1 Activation Quantization
Activation quantization reduces the precision of intermediate activation values in neural networks by converting them into low-bitwidth representations [8]. This technique is often used alongside weight quantization to further compress the model. Among various activation quantization methods, linear quantization is the most widely used. It uses a normalized scaling factor to adjust the activation values, and subsequently converts them into integers within a fixed bit range [24, 15]. The overall process typically involves three stages: forward quantization, scaling parameter computation, and dequantization [17].
3.1.1 Forward Quantization
The forward quantization process involves mapping continuous floating-point values to an integer space. This process utilizes the quantization operator $Q$, and the equation is defined as follows:
@@ 3.1.1 Forward Quantization
[…]
3.1.2 Scaling Parameter Computation
In activation quantization, setting the scaling factor $s$ is crucial for mapping various activation distributions to a fixed integer range. In this study, we employ the asymmetric quantization method, which is the most commonly used for activation quantization [29]. We calculate the scaling factor $s$ and the zero-point $z$ as follows:
@@ 3.1.2 Scaling Parameter Computation
Here, $x_{\max}$, $x_{\min}$ and $b$ represent the maximum and minimum values of the activation, and the number of quantization bits, respectively. The zero-point $z$ represents the offset used to adjust activation values. $x_{\max}$ and $x_{\min}$ determine the range of the activation distribution, which directly affects the computation of scaling factor $s$. Specifically, when $x_{\max}-x_{\min}$ is large, the scaling factor $s$ increases. This leads to a larger quantization step size, making it harder to preserve fine-grained activations. Conversely, when $x_{\max}-x_{\min}$ is small, the scaling factor $s$ also decreases. This allows for finer value distinctions and enhances the quality of representation.
3.1.3 Dequantization
The dequantization process converts quantized integer values back into their real-valued representation, as shown in the following equation:
@@ 3.1.3 Dequantization
In this process, the zero-point $z$ is subtracted to reverse the forward quantization, and the result is multiplied by the scaling factor $s$ to obtain the final value $x_{deq}$. The dequantization process restores quantized activation values to their original real-valued range. It is essential during model inference for converting final predictions into real values. This process is essential for converting the final prediction into real values during model inference.
Figure 3: Visualization of activation magnitudes for the first image in the first batch using ResNet-20 on CIFAR-10. The left column illustrates activations from the FP model, whereas the right column presents activations after layer-wise quantization. Significant distortions are observed in certain layers, particularly in the final layer.
@@ 3.2 Problem Definition
Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost. However, fixed scaling factors struggle to handle varying activations. This issue becomes more severe under low-bitwidth settings in ZSQ environments. The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
(Problem 1) Coarse-grained activation representation
In conventional activation quantization approaches, the scaling factor $s$ is fixed per layer or channel, mapping all activations to the same range. However, in low-bitwidth settings, this approach struggles to capture fine-grained activation variations from generated data.
@@
During quantization, certain layers induce severe quantization errors, leading to significant activation distortions. These distortions substantially differ from the original distribution and can directly impact the performance of the model.
@@ 4.1 Observation
ResNet-20 Layer-Q Channel-Q GranQ
@@
Table 2: Comparative analysis of averaged cosine similarity and relative error over all layers and batches. Higher cosine similarity ($\uparrow$) indicates better retention of activation information, whereas lower relative error ($\downarrow$) signifies reduced quantization distortion. Layer-Q and Channel-Q denote layer- and channel-wise quantization, respectively, whereas GranQ integrates both for improved quantization.
Layer-wise quantization effectively captures the activation distribution of each layer. However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis. Particularly in ZSQ, as shown in the FP graph in Figure 2(a), the activation distributions can differ substantially across channels even within the same layer. This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult. It consistently appears across all layers. We analyzed this phenomenon occurring across all layers and provided a summary in Table 2. In Table 2, we present the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$. From Table 2, we identify that both layer- and channel-wise quantization fail to preserve the original activation effectively. This indicates that layer- and channel-wise quantization, which independently process activations at the layer or channel level, are not suitable for ZSQ. In contrast, GranQ, which considers both layers and channels simultaneously, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error compared with those of layer-wise quantization, demonstrating its effectiveness in preserving activations.
Furthermore, Figure 3 provides a visual representation of the effect of quantization on activation values across layers. In layer-quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 3 is provided in the supplementary material, in Figures 6 and 7.
4.2 Granular Quantization
Based on our analysis, we propose GranQ (granular layer-channel quantization), a fine-grained ZSQ approach designed to minimize computational overhead, as shown in Figure 4. GranQ combines layer and channel information using layer-channel scaling and improves efficiency through vectorized quantization.
4.2.1 Layer-Channel Scaling
We propose layer-channel scaling, which enables GranQ to assign individual scaling factors to each channel, enhancing quantization precision. This approach reduces the impact of outlier activations in specific channels on the quantization process. The layer-channel scaling is defined as follows:
$\vec{x}_{\min}$ $=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$ (4)
where $\text{where }x_{l}\in\mathbb{R}^{C\times H\times W}$

### v4

# GranQ: Granular Zero-Shot Quantization with Channel-Wise Activation Scaling in QAT

@@ Abstract
Zero-shot quantization (ZSQ) enables neural network compression without original training data, making it a promising solution for restricted data access scenarios. To compensate for the lack of data, recent ZSQ methods typically rely on synthetic inputs generated from the full-precision model. However, these synthetic inputs often lead to activation distortion, especially under low-bit settings. As a result, existing methods struggle to mitigate this issue due to coarse activation scaling. To address this issue, we propose GranQ, a novel activation quantization framework that efficiently applies per-channel scaling through vectorized computation. In contrast to conventional channel-wise methods, which apply vectorization only to the quantization step, GranQ improves efficiency by vectorizing the scaling operation. This design allows GranQ to maintain fine-grained quantization granularity with minimal computational overhead, even in low-bit environments. Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet. In particular, our method achieves up to 5.45% higher accuracy in the 3-bit setting on CIFAR-100 and even surpasses the full-precision baseline on CIFAR-10. Furthermore, GranQ achieves significant speedup in quantization latency over conventional per-channel methods, demonstrating improved efficiency. With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning.
@@ 1 Introduction
Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression include quantization [1, 2, 3], pruning [4, 5, 6, 7], knowledge distillation [8, 9, 10], and neural architecture search [11, 12], as surveyed in [13, 14]. Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model. However, it often requires fine-tuning or calibration to match full-precision (FP) model performance [3]. To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data. Since the introduction of ZeroQ [15], studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
First, data generation studies focus on generating high-quality data to effectively train the Q model [16, 17, 18]. Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [19, 20]. Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [21, 22, 23, 24].
However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved. To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning. Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods. Notably, we found that layer-wise (per-tensor) quantization is no longer suitable for preserving activations in ZSQ, as it leads to coarse and inaccurate representations.
Based on this analysis, we introduce GranQ, a novel ZSQ approach that enables channel-wise activation quantization by adapting the scale to each channel of every input activation. This allows us to overcome the limitations of conventional activation quantization. This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, We apply vectorization to the scaling operation, which is not applied in conventional channel-wise quantization, reducing latency while enabling fine-grained activation quantization. Our contributions can be summarized as follows:
(a) Layer-wise Quantization
(b) GranQ (Channel-wise, Ours)
Figure 1: Comparison between (a) layer-wise quantization and (b) GranQ on the CIFAR-10. Each subfigure visualizes the 32-bit FP (left) and 3-bit quantized (right) activations of the first ResNet-20 layer. GranQ better preserves the original activation with minimal distortion.
  • We identify critical limitations of layer-wise activation quantization in low-bit ZSQ. Our findings reveal that conventional activation quantization methods relying on layer-wise granularity suffer from significant activation loss. This limitation becomes more severe in ZSQ settings with synthetic data.
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. We introduce a fine-grained activation quantization mechanism that considers every channel to reduce activation loss. Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost. We address this by applying vectorization to the scaling operation, enabling efficient per-channel quantization with minimal overhead. To the best of our knowledge, our approach is the first to address the ZSQ problem.
2 Related Work
2.1 Quantization
Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage. It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [3, 25]. PTQ applies quantization after training without further updates, which makes it efficient and easy to implement. However, it is sensitive to scaling errors and often relies on calibration to improve accuracy. In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator [3, 26]. Both methods require data—PTQ for calibration and QAT for fine-tuning. In practice, access to training data is often limited. To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data [16].
@@ 2.2 Zero-shot Quantization
As summarized in Table 1, ZSQ was initially introduced by ZeroQ [15], leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data [31, 18] and developing more effective methods to utilize them [21, 23, 24]. However, current ZSQ methods heavily rely on augmented synthetic inputs. This causes large variation in activation scales across channels. In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost. Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings [35], or exclusively in PTQ scenarios with access to a small amount of data [36, 37]. This highlights the need for efficient channel-wise activation quantization for QAT in ZSQ.
@@ 3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the appendix.
@@ 3.1 Activation Quantization
Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers [38]. It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted. This process typically consists of quantization, scaling parameter computation, and dequantization [1, 3, 39].
The quantization operator $\mathcal{Q}$ maps a floating-point value $x$ into an integer $x_{q}$ using a scaling factor $s$ and zero-point $z$:
@@
Here, $s$ normalizes the activation range, and $z$ shifts the scaled value to account for asymmetric quantization. These parameters are computed as:
@@
where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width. A large range yields a large $s$, increasing quantization error; a smaller range allows finer resolution.
During inference, the dequantization step restores quantized values to real numbers:
@@
This enables recovery of activation information needed for final predictions.
@@ 3.2 Problem Definition
Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost. However, a fixed scaling factor struggles to handle varying activations. This issue becomes more severe under low-bitwidth settings in ZSQ environments. The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
(Problem 1) Coarse Activation Quantization by Single-Range Scaling
Conventional activation quantization uses a single scaling range per layer for efficiency, which was acceptable at higher bitwidths or when real data is available. However, this design becomes problematic in ZSQ, where range adjustment relies on synthetic data. These synthetic distributions are often biased or irregular, making precise quantization more difficult.
@@
During activation quantization, certain layers induce severe quantization errors, leading to significant activation distortions. These distortions substantially differ from the original distribution and can directly impact the performance of the model.
@@ 4 Observation and Methodology
Figure 2: Activation magnitudes of ResNet-20 on CIFAR-10 using 32-bit FP (left) and 3-bit layer-wise quantization (right). Severe distortions appear, especially in the final layer.
@@ 4.1 Observation
Layer-wise quantization effectively captures the activation distribution of each layer. However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis. Particularly in ZSQ, as shown in the FP graph in Figure 1(a), the activation distributions can differ substantially across channels even within the same layer. This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult. It consistently appears across all layers. We analyzed this phenomenon occurring across all layers and provided a summary in Table 2. It presents the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$. From the results, we observe that layer-wise quantization fails to preserve activation information effectively. This reveals that using a single scaling range for activation quantization has a critical limitation, particularly in ZSQ. In contrast, GranQ, which applies scaling per channel, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error than layer-wise quantization, demonstrating its effectiveness in preserving activation information.
Furthermore, Figure 2 provides a visual representation of the effect of quantization on activation values across layers. In layer-wise quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 2 is provided in Figure 5 and in Appendix (Figure 7).
Activaiton Quantization Layer-wise Channel-wise (GranQ)
@@
Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100). GranQ (channel-wise) outperforms layer-wise quantization by achieving less distortion in activation representation, as indicated by higher similarity and lower error.
4.2 Methodology
Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead (Figure 3). GranQ applies channel-wise scaling through activation reshaping and decomposition, and achieves fast parallel processing via vectorization of scaling parameters, which enables efficient computation of both scaling and quantization.
4.2.1 Activation Reshaping and Decomposition
We propose activation reshaping and decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$. This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently. It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization. The activation reshaping and decomposition is defined as follows:
$\begin{gathered}A_{l}=\left[A_{l}(1,:,:),\,A_{l}(2,:,:),\,\dots,\,A_{l}(C,:,:)\right]\in\mathbb{R}^{C\times H\times W}\\ \vec{A}_{\min}=\left[\min_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\\ \vec{A}_{\max}=\left[\max_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\end{gathered}$ (4)
@@ 4.2.1 Activation Reshaping and Decomposition
4.2.2 Vectorization of Scaling Parameters
In the vectorization of scaling parameters stage, multiple scaling ranges are stored in a vectorized form for each channel. This allows dynamic adaptation to the diverse activation distributions induced by synthetic inputs. Thus, GranQ contrasts with traditional per-channel quantization, which uses fixed scaling parameters for each channel regardless of input variation. As shown in Equation 5, the scaling factors $\vec{s}$ and zero-points $\vec{z}$ are computed in vectorized form. The activation values are then quantized and dequantized according to Equation 6, with all operations applied in parallel across channels.
$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$ (5)
$A_{q,l}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil,A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$ (6)
This process allows efficient and adaptive quantization of activation tensors exhibiting high variation across inputs, leading to improved computational efficiency (Table 7 in appendix).
[…]
5 Experiments
5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [40], and ImageNet (ILSVRC 2012) [41] validation datasets. For the CIFAR datasets, ResNet-20 [42] was used as the quantization model, whereas ResNet-18 [42], ResNet-50 [42], and MobileNetV2 [43] were employed for ImageNet. All experiments were conducted using the SGD optimizer [44] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [16, 21, 17, 33, 34, 22, 18, 23, 24]. For data generation, we followed the AdaDFQ [17] approach based on ACGAN [45]. Further implementation details are provided in the appendix, Section A.3.2 and A.3.3. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.